# No-prior Bayesian inference reIMagined: probabilistic approximations of inferential models

### v1

@@ 5 Computation
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, (2025) developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\overline{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, (2025) can be reIMagined here as an implementation the new no-prior Bayes solution.
The strategy in Martin, (2025) is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM,” so I’ll keep it simple. For $\xi$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
@@
Of course, this vector would generally depend on $(x,\alpha)$, so I’ll denote this as $\xi(x,\alpha)$. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\xi}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\xi$. The data-driven algorithm recommended in Martin, (2025), based largely on the steps made in Cella and Martin, (2024), for selecting $\xi(x,\alpha)$ is presented in Appendix E. Once this $\xi(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
@@ 6 Example: Behrens–Fisher problem
To carry this comparison a bit further, I run a small simulation study. I focus on a rather severely imbalanced case—with $n_{1}=3$ and $n_{2}=20$—to ensure that differences in how the methods performance are apparent. Otherwise, the simulation settings are straightforward: $\Theta_{11}=2$, $\Theta_{21}=0$, $\Theta_{12}^{2}=1$, and $\Theta_{22}^{2}=2$. I generated 10000 samples according to this scheme, and Table 1 summarizes the coverage probability and expected length of the various 90% confidence intervals for $\Phi$. Notably, only the marginal possibilistic IM’s inner probabilistic approximation is able to hit the target coverage probability nearly on the nose, and, as desired, it’s more efficient than the valid-but-conservative Bayes/fiducial solution based on the right Haar prior.
@@ 7 Conclusion
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately/conservatively, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, (2025), but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from my computationally proficient readers.
@@ Acknowledgments
The above conditions are rather mild, and hold in a broad range of problems, including exponential families and many more. One further condition is required for the possibilistic Bernstein–von Mises theorem in Section 4.2, namely, that the maximum likelihood estimator is consistent, i.e., $\hat{\theta}_{X^{n}}\to\Theta$ in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$. This, of course, is not automatic, but holds quite broadly when the dimension of the parameter space is fixed, as I’m assuming here.

### v2

@@ 5 Computation
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a sufficiently fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, (2025) developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\overline{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, (2025) can be reIMagined here as an implementation of the new no-prior Bayes solution.
The strategy in Martin, (2025) is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM.” For $\xi$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
@@
Of course, this vector would generally depend on $(x,\alpha)$, so I’ll denote this as $\xi(x,\alpha)$. A cartoon depiction of the proposed strategy is shown in Figure 7. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\xi}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\xi$. The data-driven algorithm recommended in Martin, (2025), relying on the foundations laid in Cella and Martin, (2024), for selecting $\xi(x,\alpha)$ is presented in Appendix E. Once this $\xi(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
@@
[…]
Figure 7: Approximating the IM contour’s $\alpha$-level set $C_{\alpha}(x)$ (gray) by an ellipse $C_{\alpha}^{\xi}(x)$ with a “good” choice of $\xi$ (solid) and with a “bad” choice (dashed).
@@ 6 Example: Behrens–Fisher problem
For further comparison, I run a small simulation study. I focus on a rather severely imbalanced case—with $n_{1}=3$ and $n_{2}=20$—to ensure that the differences in performance are apparent. Otherwise, the simulation settings are standard: $\Theta_{11}=2$, $\Theta_{21}=0$, $\Theta_{12}^{2}=1$, and $\Theta_{22}^{2}=2$. I generated 10000 samples under this setup, and Table 1 presents the coverage probability and expected length of the various 90% confidence intervals for $\Phi$. Notably, only the marginal possibilistic IM’s inner probabilistic approximation is able to hit the target coverage probability nearly on the nose, and, as desired, it’s more efficient than the valid-but-conservative Bayes/fiducial solution based on the right Haar prior.
@@ 7 Conclusion
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately/conservatively, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, (2025), but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from computationally savvy readers.
@@ B.1 Theory
B.2 Directional data example
Let $X=(X_{1},\ldots,X_{n})$ be the observable angles for $n$ independent spins of a roulette wheel. The model assumed here is the von Mises distribution, with density
$p_{\theta}(x)=\{2\pi I_{0}(\kappa)\}^{-1}\exp\{\kappa\cos(x-\theta)\},\quad x,\theta\in[0,2\pi).$
Here $\kappa>0$ is a known concentration parameter, and $I_{0}$ denotes the Bessel function of order 0. The unknown mean angle $\Theta$ is the parameter to be inferred from the data.
Digging into this model further, using the cosine formula
$\cos(x-\theta)=\cos x\cos\theta+\sin x\sin\theta,$
it’s easy to see that the minimal sufficient statistic for this model is the pair
$Y=(\bar{C},\bar{S})=\Bigl{(}\frac{1}{n}\sum_{i=1}^{n}\cos X_{i},\,\frac{1}{n}\sum_{i=1}^{n}\sin X_{i}\Bigr{)}.$
Relative to the posited von Mises model, $Y$ can be treated as “the data.” Since $(C_{i},S_{i})=(\cos X_{i},\sin X_{i})$ denotes the Cartesian coordinates of the point on the unit circle corresponding to the angle $X_{i}$, it’s clear that $Y=(\bar{C},\bar{S})$ is generally inside the unit circle. The more “concentrated” the original data points are around a particular angle, the closer $Y$ will be to the boundary of the circle. Convert this average position to polar coordinates:
$G=\arctan(\bar{S}/\bar{C})\quad\text{and}\quad U=(\bar{C}^{2}+\bar{S}^{2})^{1/2}.$
Here $G$ is the angle the vector $(\bar{C},\bar{S})$ makes with the horizontal axis and $U$ is the Euclidean length of $(\bar{C},\bar{S})$. It’s clear that $(G,U)$ is a bijection of the minimal sufficient statistic $Y$, so write $t(X)=(G,U)$. I’m using the notation $G$ and $U$ to align with that introduced in the previous section. That is, there is a group of transformations $\mathscr{G}$ that corresponds to rotations of points in the plane; technically, $\mathscr{G}$ is the special orthogonal group $\text{SO}(2)$ consisting of orthogonal matrices with unit determinant. Then $G$ can be interpreted as both an angle or as a rotation by that angle. Similarly, the interior of the unit circle can be partitioned into distinct concentric circles, which correspond to orbits, and $U$ determines which of these orbits $Y$ sits on.
The details here align with the general setup above, so I can proceed to carry out the fiducial (or default-prior Bayes) and IM analyses. It’s a standard result (e.g., Mardia and Jupp, 2000, Eq. 4.5.5) that the joint density for $t(X)=(G,U)$ factors as
$p_{\theta}(g,u)$ $=p_{\theta}(g\mid u)\,p(u)$
=$=\{2\pi I_{0}(\kappa u)\}^{-1}\exp\{\kappa u\cos(g-\theta)\}\,p(u),$
where $p(u)$ denotes the marginal density for $U$, which doesn’t depend on $\theta$ and isn’t relevant here. The expression for the conditional density of $G$, given $U=u$, is easily seen to be that of a von Mises distribution but with concentration parameter $\kappa u$. The influence of conditioning on $U$ makes intuitive sense because, as indicated above, the statistic $U$ acts like a measure of how “concentrated” the observed angles are around a common angle. It’s the conditional density above that drives both the fiducial and IM analyses. Indeed, the no-prior Bayes (and fiducial) density is
$q_{x}(\theta)\propto\exp\{\kappa u\cos(g-\theta)\},\quad\theta\in(0,2\pi],$
and, after simplification, the IM’s possibility contour is
$\pi_{g|u}(\theta)=\mathsf{P}\{\cos H\leq\cos(g-\theta)\mid U=u\},\quad\theta\in[0,2\pi),$
where the probability is with respect to the conditional distribution of $H:=G-\Theta$, given $U=u$, which is simply a von Mises distribution with mean angle 0 and concentration parameter $\kappa u$. This contour function is easy to evaluate numerically and to plot as in the main text. By setting a sufficiently fine grid on the interval $(0,2\pi)$ and then evaluating the contour on that grid, it’s easy to obtain the level set $C_{\alpha}(x)$ for any $\alpha$. From here it’s straightforward to sample from the inner probabilistic approximation, etc.
@@ C.1 Setup and regularity conditions
The above conditions are rather mild, and hold in a broad range of problems, including exponential families. One further condition is required for the possibilistic Bernstein–von Mises theorem in Section 4.2, namely, that the maximum likelihood estimator is consistent, i.e., $\hat{\theta}_{X^{n}}\to\Theta$ in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$. This, is not automatic, but holds quite broadly when the parameter space dimension is fixed, as I’m assuming here.
C.2 Proof of Theorem 5
The intuition behind the Bernstein–von Mises result itself, and the proof that follows, is that both the inner probabilistic approximation and the Gaussian that it’s merging with are fully determined by corresponding possibility contours. So, any difference in these distributions will imply a difference in the possibility contours, but the contours themselves are merging by Theorem 4 and, therefore, any difference in the distributions must be vanishing asymptotically. The following details make this argument rigorous.
I’ll start by justifying the above claim that any difference between $\mathsf{Q}_{X^{n}}^{\star}$ and the Gaussian approximation $\mathsf{G}_{X^{n}}$ is due to differences in the corresponding possibility contours. The first observation is that $\mathsf{Q}_{X^{n}}^{\star}$ has the characterization as a uniform mixture of uniform kernels by Theorem 1. The second observation is that, similarly, $\mathsf{G}_{X^{n}}$ is the inner probabilistic approximation of the Gaussian possibility measure with contour $\gamma_{X^{n}}$ and, therefore, the characterization result in Theorem 1 applies to $\mathsf{G}_{X^{n}}$ too. That is, if $\mathsf{K}_{X^{n},\pi}^{\alpha}$ and $\mathsf{K}_{X^{n},\gamma}^{\alpha}$ are the uniform kernels supported on the level sets $\{\theta:\pi_{X^{n}}(\theta)=\alpha\}$ and $\{\theta:\gamma_{X^{n}}(\theta)=\alpha\}$, then
[…]
Then, for any Borel measurable subset $H$ of $\mathbb{T}$,
$\bigl{|}\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr{|}\leq\int_{0}^{1}\bigl{|}\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)\bigr{|}\,d\alpha,$
so $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies that there exists an $H$ such that $|\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)|$ exceeds $\varepsilon$ on a set of $\alpha$’s having positive Lebesgue measure. This implies that there exists an open subset of $H$ and a non-empty interval $(\alpha_{1},\alpha_{2})$ such that
$\pi_{X^{n}}(\theta)\in(\alpha_{1},\alpha_{2})\quad\text{and}\quad\gamma_{X^{n}}(\theta)>\alpha_{2}\quad\text{for all $\theta$ in that subset;}$;
of course, the last inequality could be replaced by “$\gamma_{X^{n}}(\theta)<\alpha_{1}$ and the positions of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ in the above display could be reversed, but these changes don’t affect the argument that follows. What matters is that there exists a point $\theta$—depending implicitly on data $X^{n}$, on $H$, and on $\varepsilon$—such that
$|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\alpha_{2}-\alpha_{1}>0.$
That Theorem 4 implies the merging of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ uniformly means that the event in the above display has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
What’s not fully satisfactory about the above conclusion is that the positive bound “$\alpha_{2}-\alpha_{1}$” is only implicit. An important point that the above argument reveals is that any differences in how $\mathsf{Q}_{X^{n}}^{\star}$ and $\mathsf{G}_{X^{n}}$ assign probability will manifest through the contour level sets. Since these are also the sets that determine the distributions themselves, it’s without loss of generality to consider hypotheses $H$ that take the form of lower level sets of either of the defining contours. For concreteness, consider
$H=\{\theta:\pi_{X^{n}}(\theta)\leq\tau\},\quad\tau\in(0,1).$
By definition of the inner probabilistic approximation, $\mathsf{Q}_{X^{n}}^{\star}(H)=\sup_{\theta\in H}\pi_{X^{n}}(\theta)=\tau$. Moreover, since $H$ is bounded and excludes the mode of $\gamma_{X^{n}}$, it’s easy to check that $\mathsf{G}_{X^{n}}(H)$ satisfies
$\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)\leq\mathsf{G}_{X^{n}}(H)\leq\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta).$
Consequently,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\leq\tau-\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\sup_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\sup_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|,$
where the second-to-last line follows because $\pi_{X^{n}}(\theta)\equiv\tau$ on $\partial H$. Similarly,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\geq\tau-\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\inf_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\inf_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\geq-\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Putting these two inequalities together gives
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Therefore,
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon\implies\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon,$
and, by Theorem 4, the latter event has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
Putting everything together, $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies existence of a set $H$ for which $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$. This implies a non-negligible difference in the contours so, without loss of generality, the problematic $H$—whose existence is guaranteed—can be taken as a lower level set corresponding to one of the contours. When $H$ is such a lower level set, $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$ implies that $|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon$ for all $\theta$ on the boundary of $H$. Since the latter event has vanishing $\mathsf{P}_{\Theta}$-probability by Theorem 4, so must the event $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$, thus proving Theorem 5.